how many moles equal 4.9 times 10^29 atoms of nitrogen
1
Expert's answer
2019-05-09T02:21:53-0400

n = N/NA = 4.9 × 1029 / 6.02 × 1023 = 0.814 × 106 mol = 0.814 MMol